# for GitHub Issue ReSolution

### v1

@@ 3.3.1 Planning
Three types of role agents engage in the planning process: the Repository Custodian, the Manager, and the Developer. The process comprises three phases: locating code files, team building, and the kick-off meeting.
Algorithm 1 Locating
1:  Input: repository: ${\mathcal{R}}_{i}$ include file $f_{i}$, GitHub issue description: $q_{x}$, LLM: ${\mathcal{L}}$
2:  Config: filter top width: $k$, prompts: ${\mathcal{P}}$
@@
7:     $f_{h},s_{h}$leftarrow $(f_{i},{\mathcal{M}})$
8:     if $\exists$ $f_{h}$ len($s_{h}$) $<$ len$(f_{i})$ then
@@
Firstly, the repository custodian employs the BM25 algorithm [28] to rank the files in the repository based on the GitHub issue description. Subsequently, the top $k$ files are selected as potential candidates for further coding. However, as described in §2.2, this simple retrieval method can introduce irrelevant files, increasing the cost and reducing the effectiveness of subsequent code changing by LLMs. Therefore, it is necessary to filter them according the relevance. While it is feasible to directly assess the relevance between each file and the issue by LLMs, many queries are redundant as some code snippets in the previous request are duplicated, leading to unnecessary costs. Considering that applying the code change often modifies a specific part of the file rather than the whole file, we propose a method to reuse the previously requested information.
As Algorithm 1 demonstrates, when a file $f_{i}$ is compared with the issue $q_{x}$ for the first time, the LLM ${\mathcal{L}}$ compresses it into a shorter code summary $s_{i}$, where $i$ denotes the file’s version number. Upon modification on $f_{i}$ to a new version, denoted as $f_{j}$, the difference, $d_{i,j}$, is obtained via the “git diff” command. If the length of $d_{i,j}$ is less than that of $f_{j}$, indicating that the version differences are smaller than the new file itself, ${\mathcal{L}}$ can summarize the code changes as a “commit” message $m_{i,j}$. $s_{i}$ together with $m_{i,j}$ makes up the description of the newer version $f_{j}$, enabling the LLM ${\mathcal{L}}$ to determine relevance $r_{j}$ in fewer length. Based on this relevance, the custodian agent filters candidate files, allowing the manager agent to define tasks with these relevant files.
@@
In this process, the manager agent has the flexibility to “recruit” team members as the issue needs. Firstly, upon receiving the candidate files, the manager begins with analyzing the GitHub issue for the repository and breaks them into detailed file-level tasks. Specifically, for each code file $f_{i}$ in the candidate set ${\mathcal{C}}_{i}^{k}$, the project manager leverages the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{4}$ and the issue description $q_{x}$ to define the corresponding file-level task $t_{i}$. One issue can be converted to multiple tasks. These tasks, along with the associated code file, are stored in a task set ${\mathcal{T}}_{i}^{k}$. Once a task is clarified, the project manager defines the personality role $r_{i}$ of the developer by invoking LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{5}$ and the task $t_{i}$. By iterating through these candidate code files, the project manager agent ultimately designs a collection of developer agent role descriptions ${\mathcal{D}}_{i}^{k}$, thus forming the development team. The details of the team building are shown in the part of Algorithm 2.
Algorithm 2 Making the Plan
1:  Input: candidate files: ${\mathcal{C}}_{i}^{k}$, issue: $q_{x}$, LLM: ${\mathcal{L}}$,
@@
After building the team, the project manager organizes a kick-off meeting. This meeting serves multiple purposes: (1) To confirm whether the tasks assigned by the project manager are reasonable and ensure that all developers in the team can collaboratively resolve the issue $q_{x}$, and (2) To determine which developers’ tasks can be executed concurrently and which tasks have dependencies need to be sorted. The meeting takes the form of a circular speech, and the project manager is responsible for opening the speech, guiding the discussion, and summarizing the results. After the meeting, developers adjust their role descriptions ${\mathcal{D}}_{i}^{k}$ based on the discussion, and the project manager, leveraging the LLM ${\mathcal{L}}$ and the prompt ${\mathcal{P}}_{7}$, generates a main work plan $c_{main}$. This work plan is presented as code, and embedded into the main program for execution.
@@ 3.3.2 Coding
Algorithm 3 Coding Task Execution
1:  Input: file-task set: ${\mathcal{T}}_{i}^{k}$, LLM: ${\mathcal{L}}$
@@
11:        $f_{i}$, old_code_part $\leftarrow$ split($f_{i}$, $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$)
12:        new_code_part $\leftarrow {\mathcal{L}}$(($f_{i}$, $t_{i}$, old_code_part), ${\mathcal{P}}_{10}$)
13:        $f_{i}^{\prime}$leftarrow replace($f_{i}$, $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$, new_code_part)
@@
Two types of agents participate in the coding process: developers and QA engineers. As outlined in Algorithm 3, for each task $t_{i}$ in the task set ${\mathcal{T}}_{i}^{k}$ and its associated code file $f_{i}$, the developer agent generates the role description of the QA engineer $a_{i}$ by leveraging the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{8}$. Subsequently, developers collaborate with their QA engineers to execute the coding tasks. During each execution of the developer, the range of lines of code that need to be modified is firstly determined as a set of intervals $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$ where $s_{i}^{\prime}$ represents the starting line number in the $i$-th hunk, and $e_{i}^{\prime}$ is the ending line number. The determination is generated by analyzing the task content $t_{i}$ and file content $f_{i}$ using the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{9}$. These intervals split the original code file $f_{i}$ into parts to be modified, old_code_part, and parts to be retained. Developers then generate new code snippets, new_code_part, by the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{10}$. The code snippets replace old_code_part, resulting in a new version of the code file $f_{i}^{\prime}$. Utilizing Git tools, the code change $\Delta d_{i}$ for this file $f_{i}$ is generated. With the code change $\Delta d_{i}$, QA engineer produce review_comment and review_decision, by the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{11}$. If the decision, review_decision, is negative (i.e., $false$), the feedback, review_comment, prompts developers to revise the code in the next attempt. This iterative process continues until the code change meets the quality standards or reaches a predefined maximum number of iterations. After the iteration, the final version of the code change, $\Delta d$, is fixed and it represents the ultimate modification result on each file. All generated final-version code changes during this process are merged into the repository-level code change ${\mathcal{D}}$ as the issue solution.
@@ 4 Experiments and Analysis
To validate the performance of our framework, we conduct experiments and compare it against other popular Large Language Models (LLMs) to demonstrate its overall effectiveness (§4.2). In addition, to further investigate the effectiveness of our framework, we assessed its performance in two processes: planning (§4.3) and coding (§4.4).
@@ 4.1 Setup
4.1.1 Dataset and Experimental Settings
In the experiments, we employ the SWE-bench dataset as the evaluation benchmark because it is the latest dataset specifically designed for evaluating the performance of the GitHub issue resolution. SWE-bench comprises $2,294$ issues extracted from $12$ popular Python repositories, representing real software evolution requirements. It includes two context settings: an Oracle retrieval and sparse retrieval. In the former setting (w/ Oracle), the LLM receives the files to modify, whereas, in the latter (w/o Oracle), the method needs to retrieve the files for modification.
Given the observation that experimental outcomes on the $25\%$ subset of SWE-bench align with those obtained from the entire dataset [15], we opt for the same $25\%$ subset previously utilized in experiments for GPT-4 according to their materials  [^3] https://drive.google.com/drive/folders/1EnrKzGAnsb_NmZKyECGmA2DrAc8ZuJ80. Moreover, the experimental scores for the five LLMs, have been made available by them  [^4] https://openreview.net/forum?id=VTF8yNQM66&noteId=lfJF38VxJr.
Our framework is flexible to integrate various LLMs. To compare with the scores reported by SWE-bench, we select GPT-4 as the base LLM. Another reason for the selection is that GPT-4 shows remarkable performance on code generation and understanding, which has been demonstrated on benchmarks such as MBPP [2] and HumanEval [9]. Claude-2 is not chosen due to the unavailability of API access.
4.1.2 Metric
Following the SWE-bench [15], we use the applied and resolved ratio to evaluate the performance under the with-Oracle setting. Specifically, the applied ratio indicates the proportion of instances where the code change is successfully generated and can be applied to the existing code repository using Git tools, i.e.,
The recall score versus file number curve is used to measure the effectiveness of file locating for the without-Oracle setting. The recall score refers to the proportion of files that are successfully located out of all the files that require modification. The formula for calculating the file location recall score for the $i$-th instance is as follows:
4.2 RQ 3: How Effective Is Our Framework?
The comparative performance analysis between our framework and other LLMs on the same dataset is presented in Table 2. The results indicate that our framework significantly outperforms other LLMs. Notably, with a resolved ratio of $13.94\%$, our framework’s effectiveness is eight-fold that of the base model, GPT-4. This substantial increase underscores our framework’s capability to harness the potential of LLMs more effectively. Furthermore, when contrasted with the previous state-of-the-art LLM, Claude-2, our framework’s resolved ratio exceeds that benchmark by more than two-fold. This superior performance unequivocally establishes the advance of our method. As Devin [34] uses a different subset from the evaluated subset of GPT-4 mentioned in the SWE-bench paper [15], the comparison between it and ours is discussed in § 5.2.
@@ 4.2 RQ 3: How Effective Is Our Framework?
    Note that this work is evaluated on a randomly chosen 25% test set, but this subset differs from the 25% subset experimented on GPT-4 in the SWE-bench.
The ablation study is designed to simulate two scenarios: (1) Without QA (w/o QA): Considering the QA engineer agent as optional within our framework, we directly evaluate the code changes generated by the developer agent, bypassing the QA process. This scenario aims to investigate the effectiveness and necessity of QA engineer review. (2) Without hints (w/o hints): Hints refer to the textual content found in the comments section of pull requests, which are typically created prior to the first commit of the pull request. The absence of hints (w/o hints) means our framework operates without any clarifications except for the issue, despite such information being available on GitHub before the issue resolution process begins. This analysis aims to explore if the participation of humans could potentially improve the success rate of issue resolution.
Our framework demonstrates the capability to significantly enhance issue resolution, even without the QA engineer or hints. It achieves a resolved ratio of $8.71$, which is five times higher than that of the base model, GPT-4. This increase underscores the contribution of other agents in our framework to its overall performance. Furthermore, integrating cooperation with QA or hints separately can further elevate the resolved ratio by $1.92$ or $1.57$, respectively. These findings serve to underscore the value of QA engineers and the participation of humans, as demonstrated by the resolved rates achieved through their integration.
For instance, to resolve the issue  [^5] https://code.djangoproject.com/ticket/30255 from the repository Django  [^6] https://github.com/django/django/, the developer modifies four hunks in two files  [^7] https://github.com/django/django/pull/12155/files, as shown in Figure 5. Despite the availability of two files in the with-Oracle setting, our proposed framework opts for modifications in only one file, as illustrated in Figure 5. Remarkably, this simpler code change enables the repository to successfully pass all requisite test cases.
Figure 3: Case from Django (Gold).
Figure 4: Case from Django (Ours).
Figure 5: Comparison of Recall Scores between MAGIS and BM25.
4.3 RQ 4: How Effective Is Our Planning Process?
To investigate the effectiveness of the planning process, we analyze the repository custodian and project manager agent, respectively. The performance of the repository custodian agent is observed in the recall score versus the file number curve, as shown in Figure 5 with the horizontal axis representing the average number of files and the vertical axis denoting the recall score associated with that number of files. This curve illustrates that our method consistently outperforms the BM25 baseline across different numbers of selected files, which suggests that our custodian can find as many relevant code files as possible with the lowest possible number of files.
For the project manager agent, we examined the alignment of its generated task descriptions with the reference code change. A higher correlation score indicates a better alignment and thus, a more accurate and effective planning direction.

### v2

@@ 3.2.1 Planning
Three types of role agents engage in the planning: Repository Custodian, Manager, and Developer. This process comprises three phases: locating code files, team building, and kick-off meeting.
Algorithm 1 Locating.
1:  Input: repository: ${\mathcal{R}}_{i}$ including files $\{f_{i}\}$, GitHub issue: $q_{x}$, LLM: ${\mathcal{L}}$
2:  Config: filter top width: $k$, prompts: ${\mathcal{P}}$, find the latest previous version of the file and its summary: $find$
@@
7:     $f_{h},s_{h}$leftarrow $find$ $(f_{i},{\mathcal{M}})$
8:     if $\exists$ $f_{h}$ and len($s_{h}$) $<$ len$(f_{i})$ then
@@
Firstly, the Repository Custodian employs the BM25 algorithm [46] to rank the files in the repository based on the GitHub issue description. Subsequently, the top $k$ files are selected as potential candidates for further coding. However, as described in §2, this simple retrieval method can introduce irrelevant files, increasing the cost and reducing the effectiveness of subsequent coding process. Therefore, we filter these files based on relevance to minimize their number. While it is feasible to directly assess the relevance between each file and the issue by LLMs, queries to the LLM may contain the same code snippets as previous ones, leading to unnecessary computational costs. Considering that applying the code change often modifies a specific part of the file rather than the entire file, we propose a memory mechanism to reuse the previously queried information.
Algorithm 1 outlines the process of locating files with our designed memory ${\mathcal{M}}$. If a file $f_{i}$ is compared for the first time with an issue $q_{x}$, the LLM ${\mathcal{L}}$ with prompt ${\mathcal{P}}_{2}$ compresses it into the summary $s_{i}$, where $i$ denotes the file’s version. This summary is shorter than the code content in the file and it is stored in memory for future reuse. If the file $f_{i}$ has been previously compared, the latest previous version ($h$) of the file $f_{h}$ can be found by the script $find$. Since $f_{i}$ can be represented as the combination of $f_{h}$ and the difference between them ($\Delta d$ that be obtained via the “git diff” command), LLMs can understand $f_{i}$ by using $f_{h}$ and $\Delta d$. If the difference is small and the file $f_{i}$ is long, it is valuable to reuse the previous summary $s_{h}$ stored in memory rather than the content of $f_{i}$. Specifically, if the length of $s_{h}$ is less than that of $f_{i}$, ${\mathcal{L}}$ with prompt ${\mathcal{P}}_{1}$ can summarize the code changes $\Delta d$ as a “commit message” $m$. The combination of $s_{h}$ and $m$ forms the description of the newer version $f_{i}$, enabling the LLM ${\mathcal{L}}$ with prompt ${\mathcal{P}}_{3}$ to determine whether it is relevant to the issue in fewer context length. Based on their relevance, the custodian agent filters irrelevant files, allowing the Manager agent to define tasks with remaining relevant files.
@@
In this process, the Manager agent has the flexibility to “recruit” team members as the issue needs. Firstly, upon receiving the located files, the Manager begins with analyzing the GitHub issue for the repository and breaks them into detailed file-level tasks. Specifically, for each code file $f_{i}$ in the candidate set ${\mathcal{C}}_{i}^{k}$, the Manager leverages the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{4}$ and the issue description $q_{x}$ to define the corresponding file-level task $t_{i}$. One issue can be converted to multiple tasks. These tasks, along with the associated code file, are stored in a task set ${\mathcal{T}}_{i}^{k}$.
Algorithm 2 Making the plan.
1:  Input: candidate files: ${\mathcal{C}}_{i}^{k}$, issue: $q_{x}$, LLM: ${\mathcal{L}}$
@@
Once a task is clarified, the Manager defines the personality role $r_{i}$ of the Developer by invoking LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{5}$ and the task $t_{i}$. By iterating through these candidate code files, the Manager agent ultimately designs a collection of Developer agent role descriptions ${\mathcal{D}}_{i}^{k}$, thus forming the development team. The details of the team building are shown in Algorithm 2. This approach simplifies the task for LLMs because each team member only needs to handle a sub-task rather than resolving the entire complex issue.
@@
After building the team, the Manager organizes a kick-off meeting. This meeting serves two purposes: ① To confirm whether the tasks assigned by the Manager are reasonable and ensure that all Developers in the team can collaboratively resolve the issue $q_{x}$, ② To determine which Developers’ tasks can be executed concurrently and which tasks have dependencies need to be sorted. The meeting takes the form of a circular speech: the Manager is responsible for opening the speech, guiding the discussion and summarizing the results, and the Developers provide their opinions based on previous discussions in turn. One example of the meeting can be found in Appendix B. After the meeting, Developers adjust their role descriptions ${\mathcal{D}}_{i}^{k}$ based on the discussion $recording$, and the Manager, leveraging the LLM ${\mathcal{L}}$ and the prompt ${\mathcal{P}}_{7}$, generates a main work plan $c_{main}$. This plan is presented as code, and embedded into the program for execution. The meeting makes collaboration among Developers more efficient and avoids potential conflicts.
@@ 3.2.2 Coding
Algorithm 3 Coding task execution.
1:  Input: file-task pairs set: ${\mathcal{T}}_{i}^{k}$, LLM: ${\mathcal{L}}$
@@
11:        $f_{i}$, old_part $\leftarrow$ split($f_{i}$, $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$)
12:        new_part $\leftarrow {\mathcal{L}}$(($f_{i}$, $t_{i}$, old_part), ${\mathcal{P}}_{10}$)
13:        $f_{i}^{\prime}$leftarrow replace($f_{i}$, $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$, new_part)
@@
Based on the empirical study on line locating and the complexity (§2), we transform the code change generation into the multi-step coding process that is designed to leverage the strengths of LLMs in code generation while mitigating their weaknesses in code change generation. Two types of agents participate in the coding process: Developers and QA Engineers. As outlined in Algorithm 3, for each task $t_{i}$ and its associated code file $f_{i}$ in ${\mathcal{T}}_{i}^{k}$, the Developer agent generates the role description of the QA Engineer $a_{i}$ by the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{8}$. Subsequently, Developers collaborate with their QA Engineers to execute the coding tasks. During each execution of the Developer, the range of lines of code that need to be modified is firstly determined as a set of intervals $\{[s_{i}^{\prime},e_{i}^{\prime}]\}$ where $s_{i}^{\prime}$ represents the starting line number in the $i$-th hunk, and $e_{i}^{\prime}$ is the ending line number. The determination is generated by analyzing the task content $t_{i}$ and file content $f_{i}$ using ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{9}$. These intervals split the original code file $f_{i}$ into parts to be modified (old_part) and parts to be retained. Developers then generate new code snippets, new_part, by ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{10}$. The code snippets replace old_part, resulting in a new version of the code file $f_{i}^{\prime}$. Utilizing Git tools, the code change $\Delta d_{i}$ for this file $f_{i}$ is generated. With the code change $\Delta d_{i}$, QA Engineer produce review_comment and review_decision, by the LLM ${\mathcal{L}}$ with the prompt ${\mathcal{P}}_{11}$. If the decision, review_decision, is negative (i.e., $false$), the feedback, review_comment, prompts Developers to revise the code in the next attempt. This iterative process continues until the code change meets the quality standards (i.e., review_decision is $true$) or reaches a predefined maximum number of iterations. After the iteration, the final version of the code change, $\Delta d$, is fixed, which is the ultimate modification result on each file. All generated final-version code changes during this process are merged into the repository-level code change ${\mathcal{D}}$ as the issue solution.
@@ 4.1 Setup
In the experiments, we employ the SWE-bench dataset as the evaluation benchmark because it is the latest dataset specifically designed for evaluating the performance of the GitHub issue resolution. SWE-bench comprises $2,294$ issues extracted from $12$ popular Python repositories, representing real software evolution requirements. Given the observation that experimental outcomes on the $25\%$ subset of SWE-bench align with those obtained from the entire dataset [27], we opt for the same $25\%$ subset previously utilized in experiments for GPT-4 according to their materials [13]. Moreover, the experimental scores for the five LLMs, have been made available by them [28].
Our framework is flexible to integrate various LLMs. To compare with the scores reported by SWE-bench, GPT-4 is selected as the base LLM. Another reason for the selection is that GPT-4 shows remarkable performance on code generation and understanding as demonstrated on benchmarks such as MBPP [2] and HumanEval [12]. Claude-2 is not chosen due to the unavailability of API access.
Following SWE-bench [27], the applied and resolved ratio is used to evaluate the performance under the setting with the files requiring modification provided. The applied ratio indicates the proportion of instances where the code change is successfully generated and can be applied to the code repository by Git. The resolved ratio refers to the proportion of instances where the code change is successfully applied and passes a series of tests. Additional elaboration is provided in Appendix C.
4.2 How Effective is Our Framework? (RQ 2)
The comparative performance analysis between our framework and other LLMs on the same dataset is presented in Tab. 2. The results indicate that our framework significantly outperforms other LLMs. Notably, with a resolved ratio of $13.94\%$, our framework’s effectiveness is eight-fold that of the base LLM, GPT-4. This substantial increase underscores our framework’s capability to harness the potential of LLMs more effectively. Furthermore, when contrasted with the previous state-of-the-art LLM, Claude-2, our framework’s resolved ratio exceeds that benchmark by more than two-fold. This superior performance unequivocally establishes the advance of our method.
@@ 4.2 How Effective is Our Framework? (RQ 2)
The ablation study is designed to simulate two scenarios: ① Without QA (w/o QA): Considering the QA Engineer agent as optional within our framework, we directly evaluate the code changes generated by the Developer agent, bypassing the QA process. This scenario aims to investigate the effectiveness and necessity of QA Engineer review. ② Without hints (w/o hints): Hints refer to the textual content found in the comments section of pull requests, which are typically created before the first commit of the pull request. This setting means our framework operates without any clarifications except for the issue, despite such information being available on GitHub before the issue resolution process begins. This analysis aims to explore if the participation of humans could potentially improve the success rate of issue resolution.
Our framework shows a significant improvement in issue resolution, even without QA or hints. It achieves a resolved ratio of $8.71$, which is five times higher than that of the base LLM. This increase underscores the contribution of other agents in MAGIS to its overall performance. Furthermore, integrating cooperation with QA or hints separately can further elevate the resolved ratio by $1.92$ or $1.57$, respectively. These findings underscore the value of QA Engineers and the participation of humans, as demonstrated by the resolved rates achieved through their integration.
For instance, to resolve the issue [17] from the repository Django [15], the developer modifies four hunks in two files [16], as shown in Fig. 15. Despite the availability of two provided files, our method opts for modifications in only one file, as illustrated in Figure 16. Remarkably, this simpler code change enables the repository to pass all requisite test cases.
Additional comparison can be found in Appendix D and E, and detailed case study is shown in Appendix H. Furthermore, the statistics on the generated code changes can be found in Appendix F.
4.3 How Effective is Our Planning Process? (RQ 3)
Figure 3: Comparison of recall scores between Ours and BM25.
@@ 4.3 How Effective is Our Planning Process? (RQ 3)
To investigate the effectiveness of the planning process, we analyze the Repository Custodian and Manager agent. The performance of the Repository Custodian agent is observed in the recall score versus the file number curve, as shown in Fig. 4. This curve demonstrates that our method consistently outperforms the BM25 baseline across varying numbers of selected files, indicating that our approach can identify the maximum number of relevant code files with the minimum selection.